# Recognition-Synergistic Scene Text Editing

### v2

@@ 4 Experiment
Methods Tamper-Syn2k Tamper-Scene ScenePair
@@
pix2pix 0.1450 9.18 34.15 127.21 13.26 -- -- -- -- --
SRNet[44] 0.0216 18.66 49.97 64.37 30.26 0.0561 14.08 26.66 49.22 17.84
SwapText[47] 0.0194 19.43 52.43 -- 54.83 -- -- -- --
MOSTEL[32] 0.0135 20.27 56.94 33.79 66.54 0.0519 14.46 27.45 49.19 37.69
DARLING[49] 0.0120 20.80 60.07 44.48 70.85 -- -- -- -- --
STEEM [50] 0.0122 20.83 72.10 24.67 78.80 -- -- -- -- --
TextCtrl[48] 0.0130 20.79 66.60 31.13 74.17 0.0447 14.99 37.56 43.78 84.67
Ours 0.0076 22.54 72.90 30.29 86.12 0.0267 17.35 46.09 41.37 91.80
@@
Ours (w/o $\mathcal{L}^{\text{cyc}}$) -- 61.2 61.2 66.8 53.8 46.4 44.8 55.7
@@ Acknowledgements
Appendix A Summary
Appendix B Details of MMDP
As described in Section 3, the input of MMPD can be denoted as $[\mathbf{E}^{\text{t}}_{T_{B}},\mathbf{E}^{\text{i}}_{I_{A}},\mathbf{E}^{\text{t}}_{\text{query}},\mathbf{E}^{\text{i}}_{\text{query}}]\in\mathbb{R}^{2(L+N)\times C}$, where $L$ presents the length of the text embeddings and $N$ presents the length of the flattened image embeddings. In our configuration, we set $L=32$, $N=256$ and $C=384$. Our MMPD consists of 12 transformer blocks, each of which includes a layer normalization layer, causal self-attention with 6 heads, and a fully connected layer.
@@ Appendix B Details of MMDP
Appendix C More Implementation Details
@@ D.2 Discrete Feature Representation
This can primarily be attributed to two factors: (1) The discretization of images introduces information distortion, resulting in poorer reconstruction quality compared to continuous representations. As shown in Table 9, for the given dataset, the reconstruction performance of the discrete form is inferior to that of the continuous form. (2) Continuous representations effectively mitigate the inherent decoding bias of the detokenizer. As discussed in Section 3.3, for continuous image features, reconstruction loss can be computed on the detokenized images, ensuring pixel-level accuracy in the final output. In contrast, for discrete representations, supervision can only be applied to the discretized image features decoded by the MMPD, leading to feature distortions during the detokenization process.
@@ Appendix F Visualization Examples of RS-STE
Figure 9: Editing results of different methods on curved text.

### v3

@@ 4 Experiment
Methods Tamper-Syn2k Tamper-Scene ScenePair #Params
@@
pix2pix 0.1450 9.18 34.15 127.21 13.26 - - - - - 54.4
SRNet[44] 0.0216 18.66 49.97 64.37 30.26 0.0561 14.08 26.66 49.22 17.84 18.95
SwapText[47] 0.0194 19.43 52.43 - 54.83 - - - - - -
MOSTEL[32] 0.0135 20.27 56.94 33.79 66.54 0.0519 14.46 27.45 49.19 37.69 54.0
DARLING[49] 0.0120 20.80 60.07 44.48 70.85 - - - - - 23.7
STEEM [50] 0.0122 20.83 72.10 24.67 78.80 - - - - - 69.6
TextCtrl[48] 0.0130 20.79 66.60 31.13 74.17 0.0447 14.99 37.56 43.78 84.67 1216.0
Ours 0.0076 22.54 72.90 30.29 86.12 0.0267 17.35 46.09 41.37 91.80 54.4
@@
Ours (w/o $\mathcal{L}^{\text{cyc}}$) - 61.2 61.2 66.8 53.8 46.4 44.8 55.7
@@ Acknowledgements
Appendix A Summary
Appendix B Details of MMDP
As described in Section 3, the input of MMPD can be denoted as $\mathbf{E}^{\text{t}}_{T_{B}},\mathbf{E}^{\text{i}}_{I_{A}},\mathbf{E}^{\text{t}}_{\text{query}},\mathbf{E}^{\text{i}}_{\text{query}}]\in\mathbb{R}^{2(L+N)\times C}$, where $L$ presents the length of the text embeddings and $N$ presents the length of the flattened image embeddings. In our configuration, we set $L=32$, $N=256$ and $C=384$. Our MMPD consists of 12 transformer blocks, each of which includes a layer normalization layer, causal self-attention with 6 heads, and a fully connected layer.
Appendix C More Implementation Details
@@
Figure 7: Editing results of different methods on curved text.
@@ D.2 Discrete Feature Representation
This can primarily be attributed to two factors: (1) The discretization of images introduces information distortion, resulting in poorer reconstruction quality compared to continuous representations. As shown in Table 9, for the given dataset, the reconstruction performance of the discrete form is inferior to that of the continuous form. (2) Continuous representations effectively mitigate the inherent decoding bias of the detokenizer. As discussed in Section 3, for continuous image features, reconstruction loss can be computed on the detokenized images, ensuring pixel-level accuracy in the final output. In contrast, for discrete representations, supervision can only be applied to the discretized image features decoded by the MMPD, leading to feature distortions during the detokenization process.